OpenStudy (gudden):

Challenge... Prove 3 = 4, without dividing or multiplying it by zero.... Also, without using that many variable addition thing.... You can just involve a single variable.... Let's see, who can do it.. Hint = ) You can skip a case, when there happen to be extra cases...

OpenStudy (rainbow_dashie):

nice pic

OpenStudy (gudden):

thanks, but you need to answer the question for medal @Rainbow_dashie

OpenStudy (rainbow_dashie):

ik im reading just wait

OpenStudy (gudden):

@hartnn ... wanna try ?

OpenStudy (mathmath333):

logic here (1)guy weds (+) (1)gal =3 (one kid)

OpenStudy (rainbow_dashie):

@gudden, You can't do this?

OpenStudy (rainbow_dashie):

Guys, let me say something, don't tell the answers, just help. Please.

OpenStudy (gudden):

I can @Rainbow_dashie And you can violate the rules of square root, ( that's one of the method.... and a great hint too )

OpenStudy (rainbow_dashie):

Oh ok your testing us

OpenStudy (anonymous):

\(3 =4\) \(Hence Proved\)... \(Ans.!\) :P

OpenStudy (gudden):

Not testing you people... just gave you a great question, to spend your time at.... Trust me, it's easy...

OpenStudy (gudden):

@waterineyes ,,,, Kindly show the work!

OpenStudy (anonymous):

Closely look, I showed.. :P

OpenStudy (gudden):

naah.! Show your work in at least 5 lines @ waterineyes... and it must be not 3 = 4 hence proved :p

OpenStudy (anonymous):

Okay: We have : LHS as 3 RHS as 4 And you know: LHS = RHS => 3 = 4 written 6 lines.. and not by your method too.. :P Anything else you want? :P

OpenStudy (gudden):

Yeah.... proper.. work..

OpenStudy (gudden):

I want proper work...:P

OpenStudy (anonymous):

How can you say that is not proper?? :P

OpenStudy (gudden):

cause you equated, LHS and RHS, without any reason....

OpenStudy (anonymous):

Okay I got it..

OpenStudy (gudden):

So many people are staring at this foolish question :P

OpenStudy (anonymous):

Should I pm you the answer or tell it here?

OpenStudy (gudden):

PM...

OpenStudy (anonymous):

Starting with: \(9-21 = 16-28\) \[9 - 21 + \frac{49}{4} = 16 - 28 + \frac{49}{4}\] \[3^2 - \frac{2.3.7}{2}= 4^2 - \frac{2.4.7}{2} + (\frac{7}{2})^2 +(\frac{7}{2})^2\] \[(3 - \frac{7}{2})^2 = (4 - \frac{7}{2})^2\]

OpenStudy (gudden):

exactly....^_^

OpenStudy (anonymous):

Take square root and cancel 7/2.

OpenStudy (gudden):

we skipped the fact of taking the modulus and still, the answer, looks correct to someone, who has no experience with this function...

OpenStudy (anonymous):

This is clearly a cheating from my side: http://engg.entrancecorner.com/old/discussion-forum/131-algebra/41606-how-to-prove-that-34.html

OpenStudy (anonymous):

Now, I am looking at what I did there.. :) The solution is not getting to the person who have posted it.. :P

geerky42 (geerky42):

It's impossible to prove it. Flawed proofs are not proofs.

OpenStudy (gudden):

haahhaha....